### v1

# DIRESA, a distance-preserving nonlinear dimension reduction technique based on regularized autoencoders

Abstract
In meteorology, finding similar weather patterns or analogs in historical datasets can be useful for data assimilation, forecasting, and postprocessing. In climate science, analogs in historical and climate projection data are used for attribution and impact studies. However, most of the time, those large weather and climate datasets are nearline. They must be downloaded, which takes a lot of bandwidth and disk space, before the computationally expensive search can be executed. We propose a dimension reduction technique based on autoencoder (AE) neural networks to compress those datasets and perform the search in an interpretable, compressed latent space. A distance-regularized Siamese twin autoencoder (DIRESA) architecture is designed to preserve distance in latent space while capturing the nonlinearities in the datasets. Using conceptual climate models of different complexities, we show that the latent components thus obtained provide physical insight into the dominant modes of variability in the system. Compressing datasets with DIRESA reduces the online storage and keeps the latent components uncorrelated, while the distance (ordering) preservation and reconstruction fidelity robustly outperform Principal Component Analysis (PCA) and other dimension reduction techniques such as UMAP or variational autoencoders.
1 Introduction
Today, weather and climate models and observation systems generate unprecedented volumes of data. For example, the Coupled Model Intercomparison Project phase 6 (CMIP6) represents more than 20 PB of climate model output data [li2023big], and the world’s Earth Observation data collections grow by around 100 PB per year [wilkinson2024].
These unwieldy archives of Earth system data provide enormous opportunities as well as challenges. They are usually stored in formats such as NetCDF, which has natively supported zlib-based compression since version 4.0. This compression technique does not consider the specific context of climate data, and as a result, many more bits of information are stored than are meaningful or necessary [klower2021compressing].
To allow for the full potential of this vast treasure of data to be unlocked, an approach is needed to reduce these datasets into a domain-aware and physically meaningful space. The approach presented in this work uses a Siamese twin AE with regularization terms to anchor the latent space to meaningful properties such as distance (ordering) preservation and statistical independence.
A major application that benefits from long records of reanalysis and climate model data is the search for analogs. Methods based on analogs, which rely on the retrieval of similar weather patterns, are used for forecasting [vandendool1989new], as Lorenz already proposed in 1969 [lorenz1969atmospheric], or for the calculation of the local attractor properties, such as the instantaneous dimension, a proxy for the predictability [faranda2017dynamical]. In analog data assimilation, the analogs can be searched in historical data archives [lguensat2017analog] or constructed via generative machine learning methods, such as variational AE [yang2021machine] or generative adversarial networks. Besides statistical and dynamical downscaling, higher resolution weather and climate information can be obtained using analog methods [zorita1999analog, ghilain2022large] or neural network-based techniques [rozoff2022comparison]. Finally, analogs are a cornerstone of a novel climate change attribution approach that accounts for changes in atmospheric circulation patterns [faranda2022climate].
As a downside to the growing datasets, however, looking up analogs in these vast records becomes increasingly computationally expensive as the dimensionality of the data and length of the dataset increases. As looking up analogs by minimizing Euclidian distance can be practically infeasible, many approaches have been proposed to quickly retrieve similar patterns of one or more atmospheric variables. These are based on dimensionality reduction (DR) techniques such as the popular classical method PCA (Principal Component Analysis). To capture the nonlinearity of the data, other techniques were developed [van2009dimensionality]. UMAP (Uniform Manifold Approximation and Projection) is one such technique [mcinnes2018umap], which can be combined with advanced search algorithms to find the analog in a fast and accurate way [franch2019mass]. DR is also combined with other data analysis techniques, such as clustering [neal2016flexible] [chattopadhyay2020predicting], or topological techniques, such as persistent homology, for finding weather regimes.
Autoencoders are another technique, based on deep learning, to produce a lower-dimensional latent representation [manning2018pca]. However, they lack the statistical independence of the latent space components and the ordering of these components. The latter can be overcome by adding a MaskLayer [royen2021masklayer]. Decorrelation of the components can be achieved by adding a covariance regularization term on the latent space layer [ladjal2019pca].
When using autoencoders, the Euclidian distance between two samples in the original space is not necessarily correlated with the distance in latent space. Our architecture uses a Siamese twin AE (see Figure 1) to preserve distance (ordering) in latent space. The two encoder branches share the weights and produce two different latent space representations. These are used for the distance loss term, similar to the invariance term used in [bardes2021vicreg], which ensures that the distance between samples in the dataset is reflected (preserved or correlated) in the distance between latent representations of those samples.
Figure 1: DIRESA: Distance regularized siamese twin AE
To better understand how the latent space relates to the dynamic variables, we demonstrate our method for a hierarchy of conceptual models, namely the well-known Lorenz ’63 model and the MAOOAM coupled ocean-atmosphere model, before applying it to reanalysis and state-of-the-art climate model data. Starting with low-dimensional, well-understood models allows us to visualize the latent space and investigate the impact of the various terms of the loss function.
2 Data
2.1 Lorenz ’63
The Lorenz ’63 system [lorenz1963deterministic] is a 3-dimensional simplified model for atmospheric convection. This system of three ordinary differential equations (ODE) displays chaos and features a strange attractor with the famous butterfly-wings shape for specific ranges of its parameters [palmer1993extended]. The 3-dimensional Lorenz butterfly is compressed into a 2-dimensional latent space by the different DR techniques. The system of ODEs with parameters $\sigma=10$, $r=28$, and $b=8/3$ was integrated using the 4th-order Runge-Kutta numerical scheme, from starting point (1,0,1) with a timestep of 0.0025. The first 1000 points are ignored as they are considered part of the transient. The next 100000 steps are split into a training dataset (first 80000), a validation (next 10000), and a test dataset (last 10000). Before feeding the data into the DR methods, it is scaled to $[0,1]$. Important to remark is that the Hausdorff fractal dimension of the Lorenz butterfly is about 2.06, so it is theoretically impossible to map the 3-dimensional dataset into two dimensions while retaining all the information.
2.2 Modular Arbitrary-Order Ocean-Atmosphere Model
MAOOAM (Modular Arbitrary-Order Ocean-Atmosphere Model) is a quasigeostrophic coupled ocean-atmosphere model for midlatitudes [de2016modular, VDDG2015]. It features a 2-layer atmosphere that is mechanically and thermodynamically coupled to a shallow-water ocean layer in the $\beta$-plane (linearized Coriolis) approximation. The DDV2016 setup was used, with an atmospheric and oceanic resolution of 2x–2y and 2x–4x, resulting in a 36-dimensional model, with ten wavenumbers for the atmospheric stream function and temperature and eight for the ocean stream function and temperature. The dataset was obtained by integrating the model with the 2nd-order Runge-Kutta numerical scheme (with nondimensional timestep 0.01) for $84e$ 8 timesteps, with a write-out every 100th timestep (1 timestep is $1/f_{0}$ s, with $f_{0}$ the Coriolis parameter at 45°N, which is about 2.7 hours). The first 1M points ($10^{8}$ timesteps) are ignored as part of the transient, required due to the slow oceanic dynamics. For the following 3M points, which represent 92000 years of data, the wavenumbers are converted to a grid of 8 x 6 and four channels (for the four variables) and then split up into a training (first 2.4M), a validation (next 300K), and a test dataset (last 300K).
3 Methods
3.1 PCA and KPCA
PCA linearly transforms the data into new coordinates, the principal components [pearson1901liii, hotelling1933analysis]. These components are the eigenvectors of the covariance matrix of the dataset. The amount of dataset variation in the direction of each eigenvector determines the ordering of the coordinates. As the covariance matrix is symmetric, the eigenvalues are real, and the eigenvectors are orthogonal. For dimensional reduction, the first principal components are taken as they explain most of the variance in the dataset.
In KPCA, the traditional PCA is combined with kernel methods [scholkopf1997kernel]. The original data points are mapped non-linearly on a high dimensional Hilbert space, where the PCA is applied. The so-called kernel trick is used to avoid working in this higher dimension (and so being highly computationally expensive). The calculations are done in the original space, replacing the standard inner product with a kernel function representing the inner product in the Hilbert space.
3.2 UMAP
UMAP is based on manifold learning and topological data analysis [mcinnes2018umap]. It creates a fuzzy topological representation of the data, a weighted graph, by connecting the points within a certain radius, which is calculated locally based on the distance to the nearest neighbors. A stochastic gradient descent algorithm with a cross-entropy loss function searches for a low-dimensional representation with the closest possible similar fuzzy topological structure.
An important hyperparameter is the number of neighbors used to construct the initial graph. Low values will focus more on the local details, while high values will better represent the big picture. A second hyperparameter is the minimum distance between points in the latent space. A low value will pack similar points closer together, while with a large value, similar points will be further apart and focus more on the global picture.
3.3 Autoencoders
An AE is a deep-learning ANN that consists of two parts: an encoder that compresses the data into a latent representation and a decoder that decompresses this representation into its original form [kramer1991nonlinear]. The AE is trained by minimizing the reconstruction loss, which ensures that the output of the decoder is as similar as possible to the encoder’s input. We use the MSE between the encoder’s input and the decoder’s output as reconstruction loss, which can also be seen as a measure for the unexplained variance of the latent representation. After the training and testing of the neural network are done, the original dataset can be compressed with the encoder.
The power of AEs comes from the fact that neural networks use non-linear activation functions, allowing them to capture more complex relationships than PCA, which is a linear transformation. Contrary to PCA, however, the latent representation usually lacks interpretability, and its components are not independent or ordered in importance.
In a VAE, the latent components are not single numbers but probability distributions, usually Gaussian distributions [kingma2022autoencoding]. The encoder has three output layers: one for the means, one for the variances (or logarithm of the variances) of the distributions, and one random sampling taken from those distributions that feed into the decoder submodel. A VAE is a generative method, as each sample of the latent distributions produces a different output after decoding. A VAE has two loss functions. The first one is the reconstruction loss, as in a standard AE. The second loss function contains a Kullback–Leibler divergence regularization term (KL loss), which forces the latent distributions to be standard normal distributions. The total loss function is the weighted average of both losses.
Figure 2: Requirements for the autoencoder. (a) Reconstructed data as close as possible to the original. (b) Distance in original space is related to distance in latent space. (c) Independence of latent components and ordering of the latent components in terms of importance.
3.4 DIRESA
Finally, we present the newly developed DIRESA DR technique. As one of the goals of this technique is to search for similar patterns, the distance between points in the original space should be reflected in their latent-space counterparts. Moreover, to provide insight into a system’s dynamics, the latent space should relate to a hierarchy of modes of variability, a kind of non-linear generalization of PCA. The AE should fulfil the following requirements (Figure 2):
  1. 1.
    Reconstructed dataset as close as possible to the original dataset
  2. 2.
    The distance between points in the original space is reflected in their distance in latent space
  3. 3.
    The latent components are independent of each other
  4. 4.
    The latent components are ordered in terms of importance
The DIRESA architecture (Figure 1) fulfils the above requirements. As in a standard AE, the reconstruction loss enforces the first requirement. We use the MSE between the original data, fed into the encoder, and the decoder output as reconstruction loss.
A shuffled version of the dataset is fed into the twin encoder, which shares the weights and biases with the original encoder. The random ordering of the shuffled dataset can be done beforehand on the complete dataset (and so does not change during training) or can be executed per batch during training by the data generator. We used the first option here to be independent of the batch size. A distance layer calculates the Euclidian distance $d(x,x^{\prime})$ between the original and shuffled input and between the latent components of the two twins, $d(z,z^{\prime})$. Minimizing the distance loss forces the relationship between those two. For DIRESA, several distance loss functions have been implemented. DIRESAMSE uses the MSE between $d(x,x^{\prime})$ and $d(z,z^{\prime})$, meaning that distance is preserved between original and latent space. The MSLE, which calculates the MSE between $log(d(x,x^{\prime})+1)$ and $log(d(z,z^{\prime})+1)$, is used for DIRESAMSLE, focusing more on short than long distances. DIRESACorr uses a distance Pearson correlation loss: $Loss_{Corr}=1-corr(d(x,x^{\prime}),d(z,z^{\prime}))$, forcing the correlation between original and latent space distances. The distance Pearson correlation loss is calculated per batch (one figure per batch and not one figure per sample as in an MSE loss function), so the batch size must be large enough to have an accurate figure. A last distance loss function is used in DIRESALogCorr, where the correlation loss is computed on the logarithm of the distances plus one.
The independence of the latent components is forced by the covariance loss $Loss_{Cov}=\sum_{i\neq j}cov^{2}_{i,j}(z)/(L*(L-1))$ which is the normalized squared sum of the non-diagonal terms of the covariance matrix, with $L$ the number of dimensions of the latent space. The denominator keeps the loss function independent of the latent space dimension. The covariance matrix is calculated by batch, so again, it is essential to have the batch size big enough so that the batch covariance is a good approximation of the dataset covariance.
The total loss is a weighted average of the three different loss components. For lowering the weight factor tuning effort, annealing [behrens2022non] is used for the covariance loss, meaning that the weight factor is 0 at the start and gradually increases. Annealing stops when the covariance loss reaches a point, where we consider the latent components independent. The covariance regularized AE (CRAE) is similar to DIRESA but with only the covariance loss and no distance loss (and so no twin encoder).
The last requirement, the ordering of the latent components, could be imposed by a Masklayer, but in practice, it was found to hamper the training of the neural network. So, the ordering is not imposed during training but is calculated afterwards. To achieve this, all encoded (latent) dataset elements are set to the average value except for one latent component. After repeating this process for all latent components and decoding them, the latent components can be ordered based on their variances in the decoded (original) space.
DR Recon MSE Cov Loss Dist Loss KL Loss
PCA 0.00191
KPCA 0.000269
UMAP 0.000282
AE 1.16e-05
BNAE 5.21e-05
CRAE 1.91e-05 7.95e-08
VAE 8.82e-05 5.20e-06
DIRESAMSE 0.000114 1.31e-05 0.000219
DIRESACorr 0.000125 3.24e-06 0.00125
Table 1: Model losses the Lorenz ’63 test dataset
Figure 3: Test dataset (blue), next to the decoded random uniform latent components (orange) for PCA (3(a)), KPCA (3(b)), UMAP (3(c)), AE (3(d)), BNAE (3(e)), CRAE (3(f)), VAE (3(g)), DIRESAMSE (3(h)) and DIRESACorr (3(i))
Figure 4: Encoded Lorenz test dataset for PCA (4(a)), KPCA (4(b)), UMAP (4(c)), AE (4(d)), BNAE (4(e)), CRAE (4(f)), VAE (4(g)), DIRESAMSE (4(h)) and DIRESACorr (4(i))
4 Results
4.1 Lorenz ’63
In our benchmark, we use nine different DR techniques to compress the 3-dimensional Lorenz butterfly into a 2-dimensional latent space: Principal Component Analysis (PCA), kernel PCA (KPCA), Uniform Manifold Approximation and Projection (UMAP), and six artificial neural networks (ANN): AE, AE with batch normalization on the latent space (BNAE), AE with covariance regularization on the latent space (CRAE), variational AE (VAE) and two distance regularized siamese twin AE (DIRESA).
The hyperparameters for the nine different DR techniques are listed in Table LABEL:table:ModelConfLorenz. Hyperparameter tuning is done using simulated annealing [kirkpatrick1983optimization] for KernelPCA and UMAP. More information on the hyperparameter tuning can be found in the Supporting Information. All ANN methods have an encoder submodel with a 3-dimensional input layer, two hidden layers with 40 and 20 fully connected nodes, and an output layer with two fully connected nodes, see Table LABEL:table:ModelDiresaLorenz.
4.1.1 Evaluation of latent and decoded samples
The different DR methods are evaluated quantitatively based on the loss function output of the test dataset (see Table 1). The properties of the latent space are assessed qualitatively by visualizing the decoding uniform random points in latent space (see Figure 3) and by visualizing the encoded test dataset in latent space (see Figure 4). In terms of reconstruction loss (or unexplained variance), PCA performs the worst (because of the linearity), and the latent space is decoded into a flat plane (see Figure 3(a)). A standard AE has the lowest reconstruction loss (the lowest unexplained variance). Still, the decoded random sample shown in Figure 3(d) is irregular and fails to generalize well outside the wings. BNAE, CRAE, and VAE are also doing well from a reconstruction loss point of view. Both DIRESA methods are doing better for the reconstruction loss than KPCA and UMAP, although a trade-off had to be made between the three different loss functions. The decoded latent space curves for both DIRESA variants follow the wings, and outside the wings, the curve is well generalized (see Figure 3(h) and 3(i)), which is also the case for BNAE (see Figure 3(e)) and VAE (see Figure 3(g)). Using an MSE or a Corr distance loss function makes little difference from a reconstruction point of view. Unlike the other methods, the UMAP decoded sample is more scattered (see Figure 3(c)) and does not lie in a 2-dimensional manifold.
Figure 4 shows the mapping of the Lorenz test dataset into latent space. The encoded latent components are uncorrelated for all DR methods except for the standard AE (see Figure 4(d)). As we can see in Figure 4(a) and 4(b), PCA and KPCA order the latent components by variance, which is not the case for the different ANNs, see Figure 4(g) for the VAE and Figure 4(h) for DIRESA, where the wing is rotated. The wing structure is well preserved with PCA, KPCA, VAE, and DIRESA methods, which is less the case for UMAP (see Figure 4(c)) and the AE. With a standard AE, no regularization is done on latent space, so we get different irregular pictures for other training runs. For BNAE, both latent components have a standard deviation of 1, so the wing structure is resized in one direction (see Figure 4(e)). CRAE has no distance regularization, so one wing is more prominent here than the other (see Figure 4(f)).
4.1.2 Distance ordering preservation
DR Corr LogCorr Can50 Pear50 Spear50 Ken50
PCA 0.997 0.996 0.117 0.839 0.889 0.834
KernelPCA 0.988 0.989 0.201 0.775 0.823 0.709
UMAP 0.918 0.892 0.300 0.711 0.751 0.603
AE 0.924 0.936 0.257 0.805 0.800 0.649
BNAE 0.987 0.987 0.155 0.853 0.874 0.769
CRAE 0.945 0.946 0.178 0.859 0.866 0.741
VAE 0.979 0.973 0.151 0.858 0.874 0.768
DIRESAMSE 0.997 0.997 0.107 0.849 0.896 0.848
DIRESACorr 0.997 0.997 0.112 0.845 0.894 0.843
Table 2: Mean distance ordering preservation between original and latent space for the Lorenz ’63 test dataset. The Corr and LogCorr columns show the Pearson correlation for Euclidean and logarithmic distances. The Can50 column gives the Canberra stability indication (smaller is better), the Pear50, Spear50 and Ken50 show the Pearson, Spearman, and Kendall correlation with a location parameter of 50. The 3 best values are shaded in red; the best is dark red.
Table 2 shows how distance ordering is preserved between original and latent space for the different DR methods. We calculated the distances from all points of the test dataset to all other points and did the same for the corresponding points in latent space. The Corr column shows the average Pearson correlation between the Euclidean distances in original and latent space. The LogCorr shows the same but for the logarithm of the distances. When searching for analogs, we are only interested in the closest distances. In the right four columns of Table 2, a location parameter 50 has been set, meaning that only the 50 closest distances (in latent space) are considered. The Can50 shows the mean Canberra stability indicator between the Euclidean distances. The Canberra stability indicator measures the difference between 2 ranked lists (smaller is better). A Canberra stability indicator of about 0 means the lists are ranked very similarly, while two random lists will result in a figure of around 1.42. Rank differences at the top of the list are more penalized than differences in the bottom part. The Pear50, Spear50 and Ken50 columns are showing Pearson’s r, Spearman’s $\rho$ and Kendall’s $\tau$ correlation with a location parameter of 50 (in the Supporting Information Table LABEL:table:LorenzDistPreservationMean100 shows the same indices for a location parameter of 100). Table 3 shows the median of the same indicators. The three best scores are shown in red, and the best score is in dark red. The Supporting Information Table LABEL:table:LorenzDistPreservationStdErr50 shows the standard error of the mean for the different indicators. Table LABEL:table:LorenzDistPreservationPvalue shows the 2 sample t-test p-value for the mean distance between DIRESA and PCA. Figure 5 shows the scatter plots of the 50 closest distances (in latent space) for the different methods for 200 random samples of the test dataset. The points in the two wings’ cross-section are where the distance ordering preservation is worst when mapped on a 2-dimensional curve.
DR Corr LogCorr Can50 Pear50 Spear50 Ken50
PCA 0.999 0.999 0.050 0.991 0.986 0.920
KernelPCA 0.992 0.993 0.157 0.894 0.898 0.750
UMAP 0.936 0.894 0.280 0.745 0.791 0.623
AE 0.938 0.947 0.233 0.845 0.833 0.670
BNAE 0.994 0.992 0.097 0.972 0.953 0.837
CRAE 0.964 0.964 0.151 0.930 0.914 0.768
VAE 0.982 0.977 0.107 0.964 0.948 0.827
DIRESAMSE 0.999 0.999 0.040 0.995 0.990 0.935
DIRESACorr 0.999 0.999 0.045 0.994 0.989 0.928
Table 3: Median distance ordering preservation between original and latent space for the Lorenz ’63 test dataset. Columns and shading as in Table 2.
Figure 5: Scatter plots of 50 closest distances for 200 random samples of the test dataset for PCA (5(a)), KPCA (5(b)), UMAP (5(c)), AE (5(d)), BNAE (5(e)), CRAE (5(f)), VAE (5(g)), DIRESAMSE (5(h)) and DIRESACorr (5(i))
Although all methods perform better than PCA concerning the reconstruction MSE, PCA outperforms almost all for the distance ordering preservation indicators, except for the DIRESA methods (which can also be seen in the scatter plots). KPCA, although grasping the curvature, does worse than PCA for distance ordering preservation. Standard AEs perform worse than the PCA on all scores because they lack regularization on latent space. Batch normalization on latent space significantly improves all scores. Looking at the median score, the DIRESAMSE method scores best on all indicators, closely followed by DIRESACorr and PCA. For the mean scores, DIRESAMSE scores best on 5 out of 6 indicators, showing that the best way to preserve distance ordering is to preserve the distance itself.
4.2 MAOOAM
Figure 6: MAOOAM: AE Reconstruction loss in function of latent size
Figure 7: Pair plot for MAOOAM latent components for PCA (7(a))
and DIRESAMSE (7(b))
For MAOOAM, the latent space dimension has been fixed to 10, based on a simulation with a standard AE (see Figure 6). When lowering the latent space dimension below 10, the reconstruction loss increases fast, which gives us an idea of the dimension of the attractor of the dynamical system.
Figure 8: Time plot for MAOOAM latent components for PCA (8(a))
and DIRESAMSE (8(b))
For MAOOAM, we will compare DIRESA with PCA. The complete DIRESA configuration can be found in the supporting information. The encoder of the DIRESA DR methods has one convolutional layer, followed by a MaxPooling and three fully connected layers, see Table LABEL:table:ModelDiresaMA. The values for all other hyperparameters can be found in LABEL:table:ModelConfMA.
4.2.1 Evaluation of latent and decoded samples
Figure 9: MAOOAM: PCA decoded latent components
Figure 10: MAOOAM: DIRESA decoded latent components
Figure 7(b) shows the pair plots for the latent components of the test dataset for PCA and DIRESA with MSE distance loss (after sorting by explained variance). The DIRESA pair plots look very similar to those of PCA. Figure 8(b) shows the time plots for the latent components of the test dataset for PCA and DIRESA with MSE distance loss. Again, the DIRESA time plot looks very similar to the PCA’s. The first four latent components with the largest explained variance have a low-frequency pattern, while the last six latent components have a high frequency. Figure 9 shows the empirical orthogonal functions (or PCA eigenvectors) for the different variables together with the explained variance (overall and per variable). To get the decoded latent components for DIRESA, all latent components of the encoded test dataset are set to their mean except for one component, which is set firstly to the mean plus one standard deviation and secondly to the mean minus one standard deviation. Both are then decoded, and the difference is taken. The result is shown in Figure 10. Decoded components for a variable with an explained variance of less than $1.5$ are faded. Again, we notice that the decoded latent DIRESA components (with MSE distance loss) are very similar to the PCA empirical orthogonal functions, except that the explained variance is higher. The unexplained variance for PCA is 3.7%, and for DIRESA 2.7%, the difference is the biggest for the atmospheric temperature, where the unexplained variance for PCA is 9.3% and for DIRESA 4.1%. We observe a strong coupling between atmospheric and ocean variables for the low-frequency components. More than 95% of the ocean stream function variance is explained by the first two components and for the ocean temperature by the first four. There is no coupling between atmospheric and ocean variables for the high-frequency components, except for component 9, which couples the atmospheric stream function with the ocean temperature.
DR Corr LogCorr Can50 Pear50 Spear50 Ken50
PCA 0.999 0.998 0.570 0.432 0.398 0.292
DIRESAMSE 0.999 0.999 0.565 0.429 0.401 0.292
DIRESAMSLE 0.999 0.998 0.518 0.490 0.453 0.338
DIRESACorr 0.999 0.999 0.553 0.453 0.417 0.304
DIRESALogCorr 0.999 0.999 0.534 0.472 0.443 0.328
Table 4: Mean distance ordering preservation between original and latent space for 200 random samples of the MAOOAM test dataset. Columns and shading as in Table 2.
4.2.2 Distance ordering preservation
Table 4 shows the distance (ordering) preservation for PCA and DIRESA with four different distance loss functions: MSE, MSLE, correlation, and correlation of the logarithm of the distances. As it is computationally expensive to calculate distances from all points of the test dataset to all other points, we took 200 random samples from the test dataset and calculated the distance to all other points. 200 random samples are enough to distinguish the best DIRESA methods (in this case, DIRESAMSLE and DIRESALogCorr) from PCA with a good p-value (see Supporting Information Table LABEL:table:MADistPreservationPvalue). The figures are shown for the same KPIs as in paragraph 4.1.2. Looking at all the distances, near and far, PCA and DIRESA methods are performing equally for the Corr KPI, and three out of the four DIRESA methods are giving slightly better figures for the LogCorr KPI (which provides more emphasis on the short distances). Looking at the 50 closest distances, all DIRESA methods outperform PCA for Can50 and Spear50, and three out of four DIRESA methods outperform PCA for the Pear50 and Ken50 KPIs. Methods with a logarithmic distance loss function yield the best results.
4.3 Conclusion
A deep ANN DR method, called DIRESA, has been developed to capture nonlinearities while preserving distance (ordering) and producing statistically independent latent components. The architecture is based on a Siamese twin AE, with three loss functions: reconstruction, covariance, and distance. An annealing method is used to automate the otherwise time-consuming process of tuning the different weights of the loss function.
DIRESA has been compared with PCA and state-of-the-art DR methods for two conceptual models, Lorenz ’63 and MAOOAM, and significantly outperforms them in terms of distance (ordering) preservation KPIs and reconstruction fidelity. We have also shown that the latent components have a physical meaning as the dominant modes of variability in the system. DIRESA correctly identifies the major coupled modes associated with the low-frequency variability of the coupled ocean-atmosphere system.
In addition to the present study, we provide an open-source Python package, based on Tensorflow [tensorflow2015-whitepaper], to build DIRESA models with convolutional and/or dense layers with one line of code [depaepe2024diresa]. On top of that, the package allows the use of custom encoder and decoder submodels to build a DIRESA model. The DIRESA package acts as a meta-model, which can use submodels with various kinds of layers, such as attention layers, and more complicated designs, such as graph neural networks or hybrid quantum neural networks. Thanks to its extensible design, the DIRESA framework can handle more complex data types, such as three-dimensional, graph, or unstructured data.
Its flexibility and robust performance make DIRESA an indispensable new tool in weather and climate science. It has all the assets to become the method of choice to distil meaningful low-dimensional representations from the ever-increasing volumes of high-resolution climate data, for applications ranging from analog retrieval to attribution studies.
5 Data availability
The data used in this study is available on the open data portal Zenodo [depaepe2024dataset].
6 Code availability
The DIRESA package source code is available on the open data portal Zenodo [depaepe2024diresa]. Other programs and scripts are available from the corresponding author upon reasonable request. Documentation can be consulted at https://diresa-learn.readthedocs.io.
7 Acknowledgements
We thank Steven Caluwaerts for the opportunity to start working on this subject as part of the Postgraduate Studies in Weather and Climate Modeling at Ghent University. We would also like to thank Herwig De Smet for reviewing this document and Alex Deckmyn for using his code for solving the Lorenz ’63 equations. LDC acknowledges support from the Belgian Science Policy Office (BELSPO) through the FED-tWIN program (Prf-2020-017) and project B2/233/P2/PRECIP-PREDICT. GDP and LDC acknowledge support from VUB through the project SRP74: LSDS. The funders played no role in the study design, data collection, analysis and interpretation of data, or the writing of this manuscript.
8 Author contributions
GDP: Data, Conceptualization, Investigation, Analysis, DIRESA Architecture, Software Development, Methodology, Writing, Visualization, Reviewing, Editing. LDC: Conceptualization, Analysis, Methodology, Writing, Reviewing, Editing, Supervision.
9 Competing interests
All authors declare no financial or non-financial competing interests.
[…]
[title=References]